emmaj2008:

Huda evaluated the expression below. Negative 3 (8 minus 5) squared minus (negative 7) = negative 3 (3) squared minus (negative 7) = negative 3 (9) minus (negative 7) = 27 minus (negative 7) = 34. What was Huda’s error?

supie:

Is is supposed to be written like this? \[-3(8-5)^2-(-7)=-3(3)^2-(-7)=-3(9)-(-7)=27-(-7)\]

emmaj2008:

i am going to re do it

supie:

Oh k.

emmaj2008:

with the question

emmaj2008:

I can re do it

emmaj2008:

cant*

supie:

Oh ok then. Is it supposed to be more than one equations?

emmaj2008:

supie:

\[-3(8-5)^2-(-7)=\]\[−3(3^2)−(−7)=\]\[(−3)(9)−(−7)=\]\[−27−(−7)\]\[=-20\]

supie:

That's the correct work, you should be able to find the error from there.